Last week, when TinyTread financial analyst team presented their analysis of the TinyTread project, Mr. Latecks asked the team to provide more information on the 4.94% discount rate used. The team has provided the following document

Shares of Badmonth Tires Inc. are currently trading at $22, and the firm has 800 thousand shares outstanding. Toni estimates that Badmonth Tires Inc.’ beta is 0.04. Using the currently accepted industry standards of a risk-free rate of 1.7%, and a market rate of 11.2%, Toni calculates the cost of equity via CAPM as:

1.7% + 0.04 (11.2%-1.7%) = 2.08%.

Further, based on Badmonth Tires Inc.’s latest balance sheet, its long-term debt is $10 million, and the book value of common equity is $3.2 million. The firm’s debt is currently selling at 95% of par, pays an annual coupon of 9%, and has a yield-to-maturity of 12%. Badmonth Tires Inc.’ marginal tax rate is 40%.

Toni estimates Badmonth Tires Inc.’ after-tax weighted average cost of capital as 4.94% based on the following:

Assess Toni’s analysis leading to the WACC and identify any potential issues. If you identify that a value is incorrect, say why and provide your alternative value
Answer #1

Cost of equity =Risk free rate+Beta*(Market Return-Risk free rate) =1.7%+0.04*(11.2%-1.7%) =2.08%
Market Value of equity =Number of Shares*Share Price =800,000*22 =17,600,000

Market Value of Debt =10000000
Cost of Debt =12%
WACC =Weight of Equity*Cost of Equity+Weight of Debt*Cost of Debt*(1-Tax Rate) =17600000/27600000*2.08%+10000000/27600000*12%*(1-40%) =3.94%